Question 12 Tries remaining: 2 Points out of 7.70 S Tax $40 P Flag question $30 $10 2 4 6 8 Calculate the Deadweight Loss that this per unit tax would generate. (Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) Answer
Answer #1

Initially the equilibrium point is where the supply curve and the demand curve Intersects each other.

Here in the figure, the equilibrium price is $20 and $6 is the equilibrium quantity.

After the tax is imposed, the supply curve shifts to the point S'.

Now we obtain a shaded region after the tax is imposed as shown in the figure below.

シMC. 30 - 爭10 lo 2Here we see that the marginal cost is not equal to the marginal benefit. Hence we obtain the deadweight loss.

The shaded region shows the deadweight loss after the tax is imposed.

To calculate it we use the formula of the area of a traingle,

Deadweight loss =1/2 * base * height.

Here base = 6 - 3 = 3, height = 15

Then the deadweight loss = 1/2 * 3 * 15 = 22.5